find n factor of h2so4 in(yes it is unbalanced equation is it necessary to make it balance?)
 
pbo2+pb+h2so4=pbso4+h2o

Grade:12

6 Answers

Naveen Kumar
9 years ago
First of all you need to understand about the balencing equation. Without balencing it you should not proceed further.
FOR CALCULATING n-FACTOR, WE SHOULD HAVE CHANGE IN OXIDATION STATE PER MOLECULE/ION, REMEMBER IT!!!!
For example if we have to get the n-factor of H2SO4, we must consider only one molecule of H2SO4.
Now balence it:
pbo2+pb+2h2so4=2pbso4+2h2o
here we can see that oxidation state of S in H2SO4 changes from +6 to +6
and so net changeper moleculeis 0.
iitian
9 years ago
0 was not in the option.
Himaja
9 years ago
The ionic reactions of the given reaction are
PbO+ SO4-2+ 4H+ + 2e-= PbSO4 + 2H2O
Pb + SO42- = PbSO4+ 2e
therefore the net reaction becomes
Pb + PbO2+ 2 h2SO4= 2PbSO4 + 2H2O
here when two moles of H2SO4 are involved 2 moles of electrons are involved in the reaction.
Hence when 1 mole of sulfuric acid is involved 1 mole of electrons are involved in the rxn
Therefore n-factor for H2SO4 = 1
Naveen Kumar
9 years ago
yes true.............
If alternatively we can see that for 2 moles of h2so4, one mole of pb is changing from 0 to +2 so for 1 h2so4, +1 is the change and so is the n-factor.
Thanks for your comments....
.......................................................
...........................................................................................................................
britto george
8 years ago
this is a simple question and you can get good things from solving this so just try
britto george
8 years ago
this is a simple question and you can get good things from solving this so just try